OpenStudy (atomicreaper):

A circle is centered at the point (5, -4) and passes through the point (-3, 2). The equation of this circle is (x +_____)2 + (y +_____)2 =______.

OpenStudy (er.mohd.amir):

with cente (h,k) and radius r equation of circle is (x-h)2+(y-k)2=r2 for r use distance formula

OpenStudy (atomicreaper):

Could you set up the equation for me? Im sure I could work it but I dont know how well I could set it up. Its been a while.

OpenStudy (er.mohd.amir):

radius=distance between point and centre=sqrt(5+3)^2+(-4-2)^2=sqrt(64+36)=sqrt(100)=10

OpenStudy (atomicreaper):

Thanks

OpenStudy (er.mohd.amir):

eq is (x-5)2+(y+4)2=100